# Causally coherent structures in turbulent dynamical systems

### v1

@@ Abstract
The magnitude of TE merits a further discussion. In the viscous wall region, an almost complete causal relationship is observed for points in the neighbourhood of $y_{A}$, see panels ($a,b$) in Figure 1. In the logarithmic region, at $y^{+}\approx 55$ (where $y_{A}$ serves as the source), the TE magnitude is approximately halved (panels $c,d$) compared to $y^{+}\approx 12$ (panels $a,b$). The reduction in TE magnitude arises from multiple factors, including interactions between the logarithmic layer and the inner and outer layers, as well as causation from alternative sources that this bi-directional approach may fail to detect. Remarkably, for $y^{+}\approx 55$ (where $y_{A}$ serves as the target), the TE magnitude increases by 15% w.r.t. the same location ($y^{+}\approx 55$), where $y_{A}$ serves as the source. Thus, in the logarithmic region, the information from the area surrounding $y_{A}$ contributes more significantly to $y_{A}$ than vice versa, consistent with the view of an independent and $Re$ invariant logarithmic layer [64, 2, 59, 9]. In the outer region ($y/\delta=0.1$), the increase in TE magnitude with $m$ reflects the slower temporal dynamics characteristic of large-scale structures. At the lowest Reynolds number, $y^{+}\approx 60$ still serves as both source and target for small $m$. Indeed, $y/\delta=0.1$ corresponds to $y^{+}\approx 77$, implying limited spatial and scale separation. In contrast, at higher $Re$, a clear outer source of information is observed. For $y/\delta=0.2,0.6,$ and $0.9$, the information flows from the outer toward the inner region. Focusing on these three $y/\delta$ locations, the filled-dotted lines (where $y_{A}$ is the target and the source is specified in the legend) consistently exhibit larger values than the circle-solid lines; see the bottom-right panel of Figure 1. As one may expect, large-scale motions at such locations, associated with the outer peak in the energy spectra at high $Re$ [26, 14, 43], have a strong causal influence on the inner layers. The analysis of such top-down causal interactions is presented in more detail later, when the net transfer entropy flux is introduced.
@@
At low Reynolds numbers, the educted CCS do not exhibit wall-normal asymmetry due to the limited separation between turbulent scales, see Figure 2. The maps in panel ($a$) are symmetric with respect to the source and target, indicating no discernible direction for transfer entropy flux at various wall-normal locations ($y^{+}=12,y^{+}\approx 55$ and $y/\delta=0.1$). The TE magnitude shows that, in the near-wall region, most of the causality originates from the surrounding area, with TE reaching its peak just a few viscous lengths away from $y_{A}$. At the highest $Re$, where the separation of scales is more pronounced, a distinct behaviour emerges. Near the wall, at $y^{+}=12$, the dominant causal interactions still originate within a few viscous lengths of the reference location. The universality of small scales is corroborated here by the symmetry of information across various Reynolds numbers. However, a weak asymmetry appears already at $y^{+}=12$, see panel $b$) in Figure 2. This asymmetry suggests an information flux from the wall to the logarithmic region. The trend is further confirmed at $y^{+}=55$, where the information flux originating from the wall and extending over a broader spanwise area, becomes more prominent. At $y^{+}=55$, the wall still imposes a causal influence, with no indication of a reverse effect. The TE magnitude decreases not only relative to that at $y^{+}=12$, but also compared to the same location at lower $Re$, see the central column in Figure 2. Such CCS resemble cross-sectional views of the well-known velocity streaks [31], which play a central role in the autonomous near-wall cycle [29]. A three-dimensional analysis is required to clarify analogies with the typical inclination angles observed in correlation maps. According to the dynamic eddy interpretation by Jiménez [30], CS are defined as strong flow features distinguished from the background via thresholding. Therefore, streaks are commonly extracted by applying thresholds to the fluctuating velocity component. Here, however, the full streamwise velocity field is considered. This typically yields irregular wall-attached layers, within which streaks appear as the very large uniform-momentum zones reported by Adrian et al. [1] and Sillero et al. [58]. Remarkably, the symmetric CCS shape changes in the outer layer ($y/\delta=0.1$) where the wall effects, i.e. the strong shear layer effects, are almost forgotten. The causality flows predominantly from the large-scale motions in the outer layer to the small-scale motions in the inner layers. The inner-layer motions are influenced by the outer layer, with top-down interactions prevailing. The causally coherent patterns extending upwards in the right panel $b$) of Figure 2 indicate such top-down interactions, in contrast to the left and central pictures. The interpretation of information-theoretic measures in fully turbulent flows, such as the ZPGTBL considered here, is considerably more complex than in simplified model flows, e.g. the Kolmogorov flow used to study the transition to turbulence [65]. At least qualitatively, the structures identified here resemble the wall-detached eddies modeled by Perry and Marusic [49] and Marusic and Perry [39]. However, further investigation is required to confirm their statistical scaling and to explain why wall-attached structures, as described by Townsend [63], are not captured by Shannon transfer entropy. The presence of such top-down causal interactions agrees with previous studies reporting the modulation of near-wall turbulence by large-scale motions [26, 46, 10, 43, 37]. Further evidence is provided below introducing the net transfer entropy flux (NTE), see Figure 3.
@@
The methodology used to extract CCS based on Shannon transfer entropy measurements is readily adaptable to different turbulent flows, or dynamical systems in general. Here, the causal metric is used to identify causally coherent patterns directly, rather than to evaluate interactions among pre-educed structures  [40], in a manner analogous to correlation maps. Two key differences distinguish CCS from classical correlation-based structures. First, CCS are typically smaller, offering a more realistic representation of the physical scale of the motions. Second, causality allows the identification of non-linear, asymmetric information flow, distinguishing sources from targets. A primary future direction involves exploring three-dimensional CCS in turbulent boundary layers, considering streamwise spatially developing structures. To address the computational cost of extracting such 3D motions, and associated with the PDF evaluation in the Shannon entropy formulation, alternative estimators may be introduced in the future, such as surrogate models capable of approximating Shannon entropy without explicitly calculating the PDFs [57]. Moreover, the choice of the source time lag ($m$) is also critical. Indeed, a 3D causally coherent structure may exhibit space and time-dependent variations with $m$.

### v2

@@ Abstract
The magnitude of TE merits a further discussion. In the viscous wall region, an almost complete causal relationship is observed for points in the neighbourhood of $y_{A}$, see left panels ($a,b$) in Figure 1. In the logarithmic region, at $y^{+}\approx 55$ (where $y_{A}$ serves as the source), the TE magnitude is approximately halved (central panels) compared to $y^{+}\approx 12$ (left panels). The reduction in TE magnitude arises from multiple factors, including interactions between the logarithmic layer and the inner and outer layers, as well as causation from alternative sources that this bi-directional approach may fail to detect. Remarkably, for $y^{+}\approx 55$ (where $y_{A}$ serves as the target), the TE magnitude increases by 15% w.r.t. the same location ($y^{+}\approx 55$), where $y_{A}$ serves as the source. Thus, in the logarithmic region, the information from the area surrounding $y_{A}$ contributes more significantly to $y_{A}$ than vice versa, consistent with the view of an independent and $Re$ invariant logarithmic layer [64, 2, 59, 9]. In the outer region ($y/\delta=0.1$), the increase in TE magnitude with $m$ reflects the slower temporal dynamics characteristic of large-scale structures. At the lowest Reynolds number, $y^{+}\approx 60$ still serves as both source and target for small $m$. Indeed, $y/\delta=0.1$ corresponds to $y^{+}\approx 77$, implying limited spatial and scale separation. In contrast, at higher $Re$, a clear outer source of information is observed. For $y/\delta=0.2,0.6,$ and $0.9$, the information flows from the outer toward the inner region. Focusing on these three $y/\delta$ locations, the filled-dotted lines (where $y_{A}$ is the target and the source is specified in the legend) consistently exhibit larger values than the circle-solid lines; see the bottom-right panel of Figure 1. As one may expect, large-scale motions at such locations, associated with the outer peak in the energy spectra at high $Re$ [26, 14, 43], have a strong causal influence on the inner layers. The analysis of such top-down causal interactions is presented in more detail later, when the net transfer entropy flux is introduced.
@@
At low Reynolds numbers, the educted CCS do not exhibit wall-normal asymmetry due to the limited separation between turbulent scales, see Figure 2. The maps in panel ($a$) are symmetric with respect to the source and target, indicating no discernible direction for transfer entropy flux at various wall-normal locations ($y^{+}=12,y^{+}\approx 55$ and $y/\delta=0.1$). The TE magnitude shows that, in the near-wall region, most of the causality originates from the surrounding area, with TE reaching its peak just a few viscous lengths away from $y_{A}$. At the highest $Re$, where the separation of scales is more pronounced, a distinct behaviour emerges. Near the wall, at $y^{+}=12$, the dominant causal interactions still originate within a few viscous lengths of the reference location. The universality of small scales is corroborated here by the symmetry of information across various Reynolds numbers. However, a weak asymmetry appears already at $y^{+}=12$, see left panel $b$) in Figure 2. This asymmetry suggests an information flux from the wall to the logarithmic region. The trend is further confirmed at $y^{+}=55$, where the information flux originating from the wall and extending over a broader spanwise area, becomes more prominent. At $y^{+}=55$, the wall still imposes a causal influence, with no indication of a reverse effect. The TE magnitude decreases not only relative to that at $y^{+}=12$, but also compared to the same location at lower $Re$, see the central column in Figure 2. Such CCS resemble cross-sectional views of the well-known velocity streaks [31], which play a central role in the autonomous near-wall cycle [29]. A three-dimensional analysis is required to clarify analogies with the typical inclination angles observed in correlation maps. According to the dynamic eddy interpretation by Jiménez [30], CS are defined as strong flow features distinguished from the background via thresholding. Therefore, streaks are commonly extracted by applying thresholds to the fluctuating velocity component. Here, however, the full streamwise velocity field is considered. This typically yields irregular wall-attached layers, within which streaks appear as the very large uniform-momentum zones reported by Adrian et al. [1] and Sillero et al. [58]. Remarkably, the symmetric CCS shape changes in the outer layer ($y/\delta=0.1$) where the wall effects, i.e. the strong shear layer effects, are almost forgotten. The causality flows predominantly from the large-scale motions in the outer layer to the small-scale motions in the inner layers. The inner-layer motions are influenced by the outer layer, with top-down interactions prevailing. The causally coherent patterns extending upwards in the right panel $b$) of Figure 2 indicate such top-down interactions, in contrast to the left and central pictures. The interpretation of information-theoretic measures in fully turbulent flows, such as the ZPGTBL considered here, is considerably more complex than in simplified model flows, e.g. the Kolmogorov flow used to study the transition to turbulence [65]. At least qualitatively, the structures identified here resemble the wall-detached eddies modeled by Perry and Marusic [49] and Marusic and Perry [39]. However, further investigation is required to confirm their statistical scaling and to explain why wall-attached structures, as described by Townsend [63], are not captured by Shannon transfer entropy. The presence of such top-down causal interactions agrees with previous studies reporting the modulation of near-wall turbulence by large-scale motions [26, 46, 10, 43, 37]. Further evidence is provided below introducing the net transfer entropy flux (NTE), see Figure 3.
@@
The methodology used to extract CCS based on Shannon transfer entropy measurements is readily adaptable to different turbulent flows, or dynamical systems in general. Here, this causal metric is used to identify causally coherent patterns directly, rather than to evaluate interactions among pre-educed structures  [40], in a manner analogous to correlation maps. Two key differences distinguish CCS from classical correlation-based structures. First, CCS are typically smaller, offering a more realistic representation of the physical scale of the motions. Second, causality allows the identification of non-linear, asymmetric information flow, distinguishing sources from targets. A primary future direction involves exploring three-dimensional CCS in turbulent boundary layers, considering streamwise spatially developing structures. To address the computational cost of extracting such 3D motions, and associated with the PDF evaluation in the Shannon entropy formulation, alternative estimators may be introduced in the future, such as surrogate models capable of approximating Shannon entropy without explicitly calculating the PDFs [57]. Moreover, the choice of the source time lag ($m$) is also critical. Indeed, a 3D causally coherent structure may exhibit space- and time-dependent variations with $m$.